OpenStudy (anonymous):

Perform the indicated operation. Add 12 feet 10 inches and 15 feet 8 inches. ___ft ____in.

OpenStudy (anonymous):

@amistre64 can u help me quick?

OpenStudy (anonymous):

12+15=27 10+8=18 27+18=45 12+8=20 10+15=25 20+25=45 12+10=22 15+8=23 22+23=45

OpenStudy (anonymous):

do u get that

OpenStudy (anonymous):

@katlin95

OpenStudy (anonymous):

yes but their is a different way to solve these